i really need help on writing "x=y-5" in standard form.
Answer by TimothyLamb(4379)

Ax + By = C
---
x = y - 5
x - y = -5
---
Answer:
(1)x + (-1)y = -5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php